next up previous
Next: About this document ...

Final Report on OD PMT Gain-Setting for SK III

K. Scholberg

Duke University

May 9, 2006


In March, I did a quick preliminary check of the gain-voltage characteristics of OD tubes newly installed in the tank, and found them roughly consistent with the ones set in the July 2005, after correcting for bad calibration of the scope used last summer (new target spe pulse height is 80 mV, as opposed to 25 mV on the bad scope channel.) However there were a few channels off by 100 V or so, and I wanted to do some further checks before Hans installed the zeners.

Portable Tube Check Setup

At the time of this check, the OD cables for the newly-installed tubes have not yet been run into the huts. The cables that go from the cable baffle to the hut for each tube are sitting on a cable tray next the the cable baffle. I used these cables in order to have the correct cable length for each tube. 1

In order to be able to check all tubes conveniently without walking in and out of the huts, I brought with me a portable Ortec 556 HV PMT supply. Next to the cable baffles, I retrieved the correct cable for each channel, plugged it into the appropriate connector at the cable baffle, and connected the other end to an old paddle card in an old crate. The Ortec HV module supplied power to this card. As in July 05, I used the paddle card scope interface to look at the tube signals on the scope (channel 1 of the infamous UCI scope, presumed good), and measured the spe pulse height as a function of HV.

Gain Check Results

I went through 103 tubes in about three days, May 6-8 2006 (I missed one one which has its cable run into hut 1,, and, which was not installed). The newly-measured tube database can be found at$\sim$schol/superk/gains/may06/tubedb_may06.txt. Tube-to-cable mapping was obtained from Bill's latest database.

Overall, results were quite consistent with those in July (after accounting for the bad scope calibration.) The recommended voltages based on the new measurements match quite well with those from July 05. New versus old recommended voltages are shown in Figure 1.

Figure 1: New (May 06) versus old (July 05) recommended HV values for newly-installed tubes. The line is $y=x$. The different colors indicate the day the tube was measured in July: blue is July 26.

However, as can be seen in the plot, there are a bunch of voltages that are off by 100-150 V with respect to last summer. I carefully rechecked several of the discrepant hut 3 tubes, but got completely consistent values each time. I think the new measurements (with the good scope channel) are solid.

It turns out that nearly all of these discrepant tubes were originally measured on a particular day- July 26- as can be seen from the colors on the plot. Blue triangles represent tubes measured on July 26: those points have a clearly different slope.

It seems likely the reason for the discrepancy is something funny with the scope: possibly channel 2 (known to be bad) flaked out and changed calibration (I actually saw some evidence of instability of that channel when I was checking it carefully in March). Another possibility is that we used a different scope channel on that day.

New recommended voltages are summarized in this file, with differences of greater than 50 V flagged.


Newly-installed OD tubes were checked with a portable setup at the cable baffle. Overall, results are consistent with those in July; however there is a batch of discrepancies with the tubes originally measured on July 26, for which I blame the flaky scope channel used in July. The new values are likely solid and should be used in place of the old ones.

next up previous
Next: About this document ...
Kate Scholberg 2006-05-09